1. A type of battery-operated led lights has a known mean lifetime 7.8 hrs with standard deviation 0.4. Assuming that the lifetimes of these led lights have a nearly symmetric/bell-curve distribution, find the percent (%) of these led lights having lifetime between 6.6 and 7.4 hours.

2. According to a medical study, it takes about 45.4 hrs, on average, for Roseola virus to fade away in infected patients (children). Assuming that the fading times of Roseola virus in infected patients have a nearly symmetric/bell-curve distribution with standard deviation 0.65, find the percent (%) of the cases where the fading time of the virus in an infected patient is more than 43.45 hours.
Answer #1

solution: - l= 7.8 6=0.4 6.6= 7.8- 3x0.4 7.8- 1x0.4 = ll - 36 = li - 6 oor 7.4 = SO, Approximately 6.6 and 7.4 115.85% of sco